In the 2010-2011 NBA regular season, the Los Angeles Lakers won 7 more than twice as many games as they lost. The Lakers played 82 games. How many wins and loses did the team have
Answer by stanbon(75887)

In the 2010-2011 NBA regular season, the Los Angeles Lakers won 7 more than twice as many games as they lost. The Lakers played 82 games. How many wins and loses did the team have
W + L = 82 games
W = 2L + 7
-------------------
Substitute for "W" and solve for "L"
2L+7 + L = 82
3L = 75
Losses = 25 games
----
Wins = 2*25+7 = 57 games
------------------
Cheers
Stan H.
----------